# Deep Learning for Retinal Degeneration Assessment: A Comprehensive Analysis of the MARIO AMD Progression Challenge

### v1

@@ Abstract
journal:  [^1] These authors contributed equally. [^2] These authors contributed equally.
@@ 1 Introduction
Age-related macular degeneration (AMD) represents one of the leading causes of irreversible visual impairment worldwide, affecting approximately 196 million people globally [1]. This progressive neurodegenerative retinal disease primarily impacts individuals over 65 years of age, causing substantial central vision loss while typically preserving peripheral vision. AMD manifests across a spectrum of severity, with advanced stages geographic atrophy (GA) and neovascular AMD (AMD) affecting approximately 20% of patients and constituting the predominant cause of severe visual impairment in developed nations [2, 3]. The multifactorial etiology of AMD, involving complex interactions between genetic susceptibility and environmental risk factors, presents significant challenges for both prevention strategies and therapeutic interventions.
The introduction of anti-vascular endothelial growth factor (anti-VEGF) therapies in 2007 revolutionized the management of AMD, demonstrating unprecedented efficacy in halting disease progression and, in some cases, restoring visual function [4, 5]. However, the success of anti-VEGF therapy is contingent upon early diagnosis, accurate disease activity assessment, and strategic treatment planning based on regular monitoring. Optical coherence tomography (OCT) has emerged as the cornerstone imaging modality in this context, providing high-resolution, three-dimensional visualization of retinal microstructure capable of detecting critical exudative markers subretinal fluid (SRF), intraretinal fluid (IRF), and intraretinal hyperreflective foci which serve as main indicators of neovascular activity (Fig.1).
@@
While artificial intelligence (AI) algorithms have demonstrated promising capabilities in identifying early and intermediate AMD stages and in predicting progression to advanced disease [6, 7], a significant gap exists in AI applications specifically designed for longitudinal monitoring of neovascular activity in patients undergoing anti-VEGF therapy. Current research has predominantly focused on cross-sectional disease classification rather than the dynamic assessment required to detect subtle temporal changes in exudative activity, information crucial for optimizing individualized treatment regimens. Developing algorithms able to accurately characterize disease evolution in these patients could substantially enhance treatment outcomes by enabling more precise, personalized anti-VEGF therapy protocols that respond directly to fluctuations in disease activity [8].
@@ 2.1 Anti-VEGF treatment paradigms
Current anti-VEGF treatment strategies for AMD follow several established paradigms, including fixed-interval dosing, pro re nata (PRN, as needed), and treat-and-extend protocols [9]. Each approach presents distinct advantages and limitations regarding treatment burden, visual outcomes, and healthcare resource utilization. The fixed-interval dosing regimen, derived from pivotal clinical trials like MARINA and ANCHOR [4, 5], involves regular monthly injections regardless of disease activity. While this approach provides optimal visual outcomes, it imposes substantial treatment burden on patients and healthcare systems. Conversely, the PRN approach involves treatment only when active disease is detected during monthly monitoring visits, reducing injection frequency but requiring frequent clinical assessment [10]. The treat-and-extend protocol, which has gained prominence in recent years, aims to optimize the treatment-monitoring balance by gradually extending the interval between treatments when disease stability is achieved, and shortening intervals when recurrent activity is detected [11]. This personalized approach necessitates accurate and consistent evaluation of disease activity markers, particularly fluid presence and evolution. This is precisely the clinical need that the MARIO challenge addresses.
@@ 2.2 Clinical significance of fluid dynamics
The presence, location, and quantity of retinal fluid serve as primary biomarkers for AMD activity and treatment response [12]. Different fluid compartments—subretinal fluid (SRF), intraretinal fluid (IRF), and sub-retinal pigment epithelium fluid—exhibit distinct associations with visual outcomes and treatment response patterns. Studies have demonstrated that persistent IRF strongly correlates with poorer visual outcomes and may indicate irreversible retinal damage [13, 14], while moderate SRF may be tolerated without significant visual impact in some patients [15]. This nuanced understanding of fluid dynamics has led to more sophisticated treatment decision-making, moving beyond binary dry/wet classification toward quantitative assessment of specific fluid patterns and their evolution over time. The development of AI algorithms able to detect subtle changes in fluid distribution and volume across sequential visits therefore addresses a critical unmet need in clinical practice: the ability to objectively quantify disease activity dynamics and personalize treatment decisions based on individual response patterns rather than on standardized protocols.
@@ 2.3.1 Registration and alignment
Accurate spatial alignment between sequential OCT volumes is fundamental for reliable detection of temporal changes. Patient movement, variations in scan protocols, and alterations in retinal morphology due to disease progression or treatment effects can complicate registration [16]. While commercial OCT systems incorporate basic eye-tracking functionality, suboptimal alignment remains common in clinical datasets. Advanced registration techniques are required to establish precise spatial correspondence across sequential scans. These approaches must account for both global transformation parameters and local deformations resulting from disease activity changes. Methods ranging from feature-based registration to deformable transformation models have been explored, though their application in routine clinical analysis remains limited [17].
@@ 2.3.2 Handling irregular time intervals
Clinical follow-up schedules in AMD management frequently involve irregular time intervals between visits, with intervals ranging from 4 weeks to several months depending on disease activity and treatment protocol. This temporal irregularity poses significant challenges for conventional sequence modeling approaches, which typically assume uniform sampling intervals. Recent deep learning approaches have adapted to address this challenge, including temporal convolutional networks with dilated convolutions [18], continuous-time models like neural ordinary differential equations (NODEs) [19], and attention-based architectures that can model long-range dependencies regardless of temporal distance [20]. The MARIO challenge provides an opportunity to evaluate these approaches in a standardized clinical context.
@@ 2.3.3 Balancing sensitivity and specificity
The clinical utility of AMD monitoring algorithms depends on achieving an optimal balance between sensitivity (detecting all instances of disease activity) and specificity (avoiding false positive detections that could lead to unnecessary treatment). This balance is particularly crucial in treatment decision-making, where false negatives could result in undertreatment and disease progression, while false positives might lead to overtreatment and increased iatrogenic risks [21]. Conventional evaluation metrics like accuracy may be insufficient or even misleading in this context, particularly when dealing with class imbalance or when considering the differential clinical impact of various error types. The MARIO challenge therefore incorporates clinically relevant evaluation metrics that reflect the actual utility of algorithms in treatment decision support.
Research in AI-assisted AMD analysis using OCT imaging has evolved substantially in recent years, focusing on several interconnected domains [22, 23, 24, 25]: automated retinal layer segmentation, detection and quantification of exudative features, longitudinal analysis of disease progression, and cross-domain generalizability. These research directions collectively provide the foundation for comprehensive understanding of AMD pathophysiology and evidence-based clinical decision-making. We review these key areas with particular emphasis on their relevance to monitoring disease progression in patients undergoing anti-VEGF treatments.
@@ 2.3.4 Exudative feature detection and quantification
The detection and precise quantification of exudative markers including SRF, IRF, and hyperreflective foci - represents a critical component in evaluating neovascular AMD activity, as these features directly inform anti-VEGF treatment decisions and retreatment intervals. Early approaches relied on traditional computer vision techniques with handcrafted feature extraction, but deep learning architectures have since transformed the landscape of exudative feature detection [26, 27]. Convolutional neural networks, particularly implementations of mask R-CNN, U-Net derivatives, and DenseNet-based architectures, have significantly improved accuracy in identifying and segmenting fluid compartments and hyperreflective material [28]. However, these models continue to face challenges with variations in image quality, scan protocols, and patient-specific anatomical differences. Recent advancements have explored specialized architectural adaptations, including attention mechanisms [29], Transformer-based models [30], and hybrid approaches that integrate anatomical priors with data-driven learning [31].
While these innovations have improved performance in controlled settings, the reliable detection of subtle fluid changes over time remains challenging, particularly in patients undergoing active treatment. The MARIO challenge format, with its focus on patients receiving ongoing anti-VEGF therapy, provides an ideal testbed for evaluating models capable of sensitive and specific fluid detection across sequential visits. This capability is crucial for longitudinal disease monitoring and capturing the temporal evolution of exudative markers, which directly impacts treatment decision-making [8, 32].
@@ 2.3.5 Longitudinal analysis and progression modeling
Self-supervised learning techniques have demonstrated promise in modeling disease trajectories and predicting the onset of advanced AMD stages without requiring extensive labeled data [33, 34]. These approaches enable the extraction of meaningful representations from temporal sequences of OCT volumes, capturing subtle changes that might escape human detection. Recurrent neural networks, particularly LSTM and GRU architectures, have been applied to model sequential dependencies in OCT data. However, they often struggle with irregular sampling intervals common in clinical practice [35].
More recently, differential equation-based models, especially NODEs, have emerged as powerful frameworks for modeling continuous-time disease dynamics in various conditions including AMD [36, 37]. These approaches offer distinct advantages in handling irregular sampling intervals and providing interpretable trajectories of disease states, allowing for more natural integration of multimodal data including imaging biomarkers and clinical factors.
@@ 2.3.6 Domain adaptation and generalizability
Cross-domain generalizability remains one of the most significant challenges in clinical OCT analysis for AMD [38, 39]. Substantial variability in OCT data can arise from differences in scanner types, acquisition protocols, image quality, and patient demographics, often resulting in significant performance degradation when algorithms are applied outside their training domain. Various domain adaptation techniques have been proposed to address these challenges, including adversarial learning approaches [40], transfer learning strategies [41], and generative models such as CycleGANs for style transfer between domains [42]. These methods aim to reduce the domain shift between source and target distributions, thereby improving model robustness across heterogeneous data sources. However, the fundamental limitation of restricted data diversity persists in most current research, with algorithms developed and validated predominantly on homogeneous populations showing diminished performance when applied to underrepresented groups.
Addressing algorithmic bias is therefore essential for ensuring equitable performance across diverse populations [43, 44]. The MARIO challenge uniquely addresses these limitations by providing a dataset inclusive of both African and European patient data, promoting the development of adaptable models suitable for global deployment in diverse clinical settings.
@@ 2.3.7 Publicly-available datasets and benchmarking efforts
Public datasets and standardized benchmarks have been instrumental in advancing AMD research, particularly in the domains of fluid detection and retinal layer segmentation. Notable contributions include the RETOUCH challenge dataset [45], which features OCT images from multiple devices and various retinal diseases, establishing important benchmarks for fluid detection and segmentation tasks. Similarly, the OCTAGON dataset [46] provides valuable resources for algorithm development and validation in retinal OCT analysis. However, these existing datasets have limitations that restrict their utility for developing comprehensive AMD monitoring solutions. Most critically, they often lack:
@@ 3 Challenge description
The MARIO challenge is designed following the BIAS Reporting Guideline [47] for enhanced quality and transparency of biomedical research. The proposal was approved after two rounds of MICCAI review followed by a call for participation circulated online and offline. The challenge was officially launched on April 1, 2024, and run through a 6-months window as shown in Fig. 3 using the codabench platform for hosting [48]. This period is characterized by several activities, such as the release of training data, customized metrics library, slack communication channel, a snippet of “getting started” code, GitHub repositories, etc., to guide and support the participants’ method development. Fig.2 presents the numbers of teams and associated countries that register for the challenge and also the countries associated to the finalists. The participating teams develop their novel methods, fine-tune a state-of-the-art method, or improve on existing solutions during this period. The challenge timeline also involves a validation phase, harnessed by the use of a self-validation system, validation data samples, a Docker template, and guidelines provided to facilitate method submission. The whole process is concluded with the presentation of the method, results, and award winners at MICCAI 2024 conference in Marrakesh on October 10, 2024.
@@ 4 Data description
The MARIO database, comprises a comprehensive longitudinal collection of ophthalmic data, facilitating detailed analysis of AMD. As illustrated in Figure 6, the dataset’s composition and characteristics provide a robust foundation for investigating disease progression and treatment efficacy. The gender distribution reveals a notable skew towards female patients, with 84 female participants compared to 52 male participants, suggesting a potential demographic bias in the patient population or a higher prevalence of AMD among females [49]. The imaging frequency followed standardized protocols for most patients, with 785 examinations performed using 19 B-scans and 337 examinations using 25 B-scans. The observed variation in B-scan count likely reflects differences in clinical assessments or adjustments to protocols tailored to individual patient needs. In terms of follow-up timing, the majority of patients demonstrated short inter-visit intervals, suggesting consistent and diligent monitoring. Nevertheless, the presence of outliers, with inter-visit periods extending to 1600 days, points to instances of irregular follow-up, possibly due to patient compliance issues or clinical scheduling limitations.
@@ 4.4 Justification for dataset size
Data export was manual and, therefore, time-consuming. In addition, manually annotating thousands of image pairs by two retinal experts was also time-consuming. This was the reason for not collecting a larger dataset. However, this remains larger than a similar study published by one of the organizers [50], which involved 70 patients, as opposed to 136 here.

### v3

# Deep Learning for Retinal Degeneration Assessment: A Comprehensive Analysis of the MARIO , Challenge

@@ 1 Introduction
Age-related macular degeneration (AMD) represents one of the leading causes of irreversible visual impairment worldwide, affecting approximately 196 million people globally [SCHULTZ20211792]. This progressive neurodegenerative retinal disease primarily impacts individuals over 65 years of age, causing substantial central vision loss while typically preserving peripheral vision. AMD manifests across a spectrum of severity, with advanced stages geographic atrophy (GA) and neovascular AMD (AMD) affecting approximately 20% of patients and constituting the predominant cause of severe visual impairment in developed nations [Wong2014, Fleckenstein2018]. The multifactorial etiology of AMD, involving complex interactions between genetic susceptibility and environmental risk factors, presents significant challenges for both prevention strategies and therapeutic interventions.
The introduction of anti-vascular endothelial growth factor (anti-VEGF) therapies in 2007 revolutionized the management of AMD, demonstrating unprecedented efficacy in halting disease progression and, in some cases, restoring visual function [Rosenfeld2006, Brown2009]. However, the success of anti-VEGF therapy is contingent upon early diagnosis, accurate disease activity assessment, and strategic treatment planning based on regular monitoring. Optical coherence tomography (OCT) has emerged as the cornerstone imaging modality in this context, providing high-resolution, three-dimensional visualization of retinal microstructure capable of detecting critical exudative markers subretinal fluid (SRF), intraretinal fluid (IRF), and intraretinal hyperreflective foci which serve as main indicators of neovascular activity (Fig.1).
@@
While artificial intelligence (AI) algorithms have demonstrated promising capabilities in identifying early and intermediate AMD stages and in predicting progression to advanced disease [Yim2020, Grassmann2018], a significant gap exists in AI applications specifically designed for longitudinal monitoring of neovascular activity in patients undergoing anti-VEGF therapy. Current research has predominantly focused on cross-sectional disease classification rather than the dynamic assessment required to detect subtle temporal changes in exudative activity, information crucial for optimizing individualized treatment regimens. Developing algorithms able to accurately characterize disease evolution in these patients could substantially enhance treatment outcomes by enabling more precise, personalized anti-VEGF therapy protocols that respond directly to fluctuations in disease activity [SchmidtErfurth2022].
@@ 2.1 Anti-VEGF treatment paradigms
Current anti-VEGF treatment strategies for AMD follow several established paradigms, including fixed-interval dosing, pro re nata (PRN, as needed), and treat-and-extend protocols [Lanzetta2021]. Each approach presents distinct advantages and limitations regarding treatment burden, visual outcomes, and healthcare resource utilization. The fixed-interval dosing regimen, derived from pivotal clinical trials like MARINA and ANCHOR [Rosenfeld2006, Brown2009], involves regular monthly injections regardless of disease activity. While this approach provides optimal visual outcomes, it imposes substantial treatment burden on patients and healthcare systems. Conversely, the PRN approach involves treatment only when active disease is detected during monthly monitoring visits, reducing injection frequency but requiring frequent clinical assessment [Lalwani2009]. The treat-and-extend protocol, which has gained prominence in recent years, aims to optimize the treatment-monitoring balance by gradually extending the interval between treatments when disease stability is achieved, and shortening intervals when recurrent activity is detected [Freund2015]. This personalized approach necessitates accurate and consistent evaluation of disease activity markers, particularly fluid presence and evolution. This is precisely the clinical need that the MARIO challenge addresses.
@@ 2.2 Clinical significance of fluid dynamics
The presence, location, and quantity of retinal fluid serve as primary biomarkers for AMD activity and treatment response [Schmidt-Erfurth2018]. Different fluid compartments—subretinal fluid (SRF), intraretinal fluid (IRF), and sub-retinal pigment epithelium fluid—exhibit distinct associations with visual outcomes and treatment response patterns. Studies have demonstrated that persistent IRF strongly correlates with poorer visual outcomes and may indicate irreversible retinal damage [Blumenkranz2010, Sharma2016], while moderate SRF may be tolerated without significant visual impact in some patients [Guymer2019]. This nuanced understanding of fluid dynamics has led to more sophisticated treatment decision-making, moving beyond binary dry/wet classification toward quantitative assessment of specific fluid patterns and their evolution over time. The development of AI algorithms able to detect subtle changes in fluid distribution and volume across sequential visits therefore addresses a critical unmet need in clinical practice: the ability to objectively quantify disease activity dynamics and personalize treatment decisions based on individual response patterns rather than on standardized protocols.
@@ 2.3.1 Registration and alignment
Accurate spatial alignment between sequential OCT volumes is fundamental for reliable detection of temporal changes. Patient movement, variations in scan protocols, and alterations in retinal morphology due to disease progression or treatment effects can complicate registration [Lang2016]. While commercial OCT systems incorporate basic eye-tracking functionality, suboptimal alignment remains common in clinical datasets. Advanced registration techniques are required to establish precise spatial correspondence across sequential scans. These approaches must account for both global transformation parameters and local deformations resulting from disease activity changes. Methods ranging from feature-based registration to deformable transformation models have been explored, though their application in routine clinical analysis remains limited [Golabbakhsh2013].
@@ 2.3.2 Handling irregular time intervals
Clinical follow-up schedules in AMD management frequently involve irregular time intervals between visits, with intervals ranging from 4 weeks to several months depending on disease activity and treatment protocol. This temporal irregularity poses significant challenges for conventional sequence modeling approaches, which typically assume uniform sampling intervals. Recent deep learning approaches have adapted to address this challenge, including temporal convolutional networks with dilated convolutions [Lea2017], continuous-time models like neural ordinary differential equations (NODEs) [Chen2018], and attention-based architectures that can model long-range dependencies regardless of temporal distance [Vaswani2017]. The MARIO challenge provides an opportunity to evaluate these approaches in a standardized clinical context.
@@ 2.3.3 Balancing sensitivity and specificity
The clinical utility of AMD monitoring algorithms depends on achieving an optimal balance between sensitivity (detecting all instances of disease activity) and specificity (avoiding false positive detections that could lead to unnecessary treatment). This balance is particularly crucial in treatment decision-making, where false negatives could result in undertreatment and disease progression, while false positives might lead to overtreatment and increased iatrogenic risks [Liu_2018Hanruo]. Conventional evaluation metrics like accuracy may be insufficient or even misleading in this context, particularly when dealing with class imbalance or when considering the differential clinical impact of various error types. The MARIO challenge therefore incorporates clinically relevant evaluation metrics that reflect the actual utility of algorithms in treatment decision support.
Research in AI-assisted AMD analysis using OCT imaging has evolved substantially in recent years, focusing on several interconnected domains [Crincoli2024, Romond2021, Muntean2023, SchmidtErfurth2016]: automated retinal layer segmentation, detection and quantification of exudative features, longitudinal analysis of disease progression, and cross-domain generalizability. These research directions collectively provide the foundation for comprehensive understanding of AMD pathophysiology and evidence-based clinical decision-making. We review these key areas with particular emphasis on their relevance to monitoring disease progression in patients undergoing anti-VEGF treatments.
@@ 2.3.4 Exudative feature detection and quantification
The detection and precise quantification of exudative markers including SRF, IRF, and hyperreflective foci - represents a critical component in evaluating neovascular AMD activity, as these features directly inform anti-VEGF treatment decisions and retreatment intervals. Early approaches relied on traditional computer vision techniques with handcrafted feature extraction, but deep learning architectures have since transformed the landscape of exudative feature detection [hassan_deep_2021, Schlegl2018]. Convolutional neural networks, particularly implementations of mask R-CNN, U-Net derivatives, and DenseNet-based architectures, have significantly improved accuracy in identifying and segmenting fluid compartments and hyperreflective material [Venhuizen2020]. However, these models continue to face challenges with variations in image quality, scan protocols, and patient-specific anatomical differences. Recent advancements have explored specialized architectural adaptations, including attention mechanisms [Xu2023], Transformer-based models [Melinak2023], and hybrid approaches that integrate anatomical priors with data-driven learning [Roy2017].
While these innovations have improved performance in controlled settings, the reliable detection of subtle fluid changes over time remains challenging, particularly in patients undergoing active treatment. The MARIO challenge format, with its focus on patients receiving ongoing anti-VEGF therapy, provides an ideal testbed for evaluating models capable of sensitive and specific fluid detection across sequential visits. This capability is crucial for longitudinal disease monitoring and capturing the temporal evolution of exudative markers, which directly impacts treatment decision-making [SchmidtErfurth2022, Simader2014].
@@ 2.3.5 Longitudinal analysis and progression modeling
Self-supervised learning techniques have demonstrated promise in modeling disease trajectories and predicting the onset of advanced AMD stages without requiring extensive labeled data [rivail2019modeling, Jung2024]. These approaches enable the extraction of meaningful representations from temporal sequences of OCT volumes, capturing subtle changes that might escape human detection. Recurrent neural networks, particularly LSTM and GRU architectures, have been applied to model sequential dependencies in OCT data. However, they often struggle with irregular sampling intervals common in clinical practice [Lad2022].
More recently, differential equation-based models, especially NODEs, have emerged as powerful frameworks for modeling continuous-time disease dynamics in various conditions including AMD [Yellapragada2022, Chakravarty2024]. These approaches offer distinct advantages in handling irregular sampling intervals and providing interpretable trajectories of disease states, allowing for more natural integration of multimodal data including imaging biomarkers and clinical factors.
@@ 2.3.6 Domain adaptation and generalizability
Cross-domain generalizability remains one of the most significant challenges in clinical OCT analysis for AMD [kugelman2022review, MATTA2024109256]. Substantial variability in OCT data can arise from differences in scanner types, acquisition protocols, image quality, and patient demographics, often resulting in significant performance degradation when algorithms are applied outside their training domain. Various domain adaptation techniques have been proposed to address these challenges, including adversarial learning approaches [guan2021domain], transfer learning strategies [liu2022deep], and generative models such as CycleGANs for style transfer between domains [Gomariz2022]. These methods aim to reduce the domain shift between source and target distributions, thereby improving model robustness across heterogeneous data sources. However, the fundamental limitation of restricted data diversity persists in most current research, with algorithms developed and validated predominantly on homogeneous populations showing diminished performance when applied to underrepresented groups.
Addressing algorithmic bias is therefore essential for ensuring equitable performance across diverse populations [ueda2024fairness, LIM2024100096]. The MARIO challenge uniquely addresses these limitations by providing a dataset inclusive of both African and European patient data, promoting the development of adaptable models suitable for global deployment in diverse clinical settings.
@@ 2.3.7 Publicly-available datasets and benchmarking efforts
Public datasets and standardized benchmarks have been instrumental in advancing AMD research, particularly in the domains of fluid detection and retinal layer segmentation. Notable contributions include the RETOUCH challenge dataset [Bogunovic2019], which features OCT images from multiple devices and various retinal diseases, establishing important benchmarks for fluid detection and segmentation tasks. Similarly, the OCTAGON dataset [Daz2019] provides valuable resources for algorithm development and validation in retinal OCT analysis. However, these existing datasets have limitations that restrict their utility for developing comprehensive AMD monitoring solutions. Most critically, they often lack:
@@ 3 Challenge description
The MARIO challenge is designed following the BIAS Reporting Guideline [MaierHein2018] for enhanced quality and transparency of biomedical research. The proposal was approved after two rounds of MICCAI review followed by a call for participation circulated online and offline. The challenge was officially launched on April 1, 2024, and run through a 6-months window as shown in Fig. 3 using the codabench platform for hosting [codabench]. This period is characterized by several activities, such as the release of training data, customized metrics library, slack communication channel, a snippet of “getting started” code, GitHub repositories, etc., to guide and support the participants’ method development. Fig.2 presents the numbers of teams and associated countries that register for the challenge and also the countries associated to the finalists. The participating teams develop their novel methods, fine-tune a state-of-the-art method, or improve on existing solutions during this period. The challenge timeline also involves a validation phase, harnessed by the use of a self-validation system, validation data samples, a Docker template, and guidelines provided to facilitate method submission. The whole process is concluded with the presentation of the method, results, and award winners at MICCAI 2024 conference in Marrakesh on October 10, 2024.
@@ 4 Data description
The MARIO database, comprises a comprehensive longitudinal collection of ophthalmic data, facilitating detailed analysis of AMD. As illustrated in Figure 6, the dataset’s composition and characteristics provide a robust foundation for investigating disease progression and treatment efficacy. The gender distribution reveals a notable skew towards female patients, with 84 female participants compared to 52 male participants, suggesting a potential demographic bias in the patient population or a higher prevalence of AMD among females [Rudnicka2012]. The imaging frequency followed standardized protocols for most patients, with 785 examinations performed using 19 B-scans and 337 examinations using 25 B-scans. The observed variation in B-scan count likely reflects differences in clinical assessments or adjustments to protocols tailored to individual patient needs. In terms of follow-up timing, the majority of patients demonstrated short inter-visit intervals, suggesting consistent and diligent monitoring. Nevertheless, the presence of outliers, with inter-visit periods extending to 1600 days, points to instances of irregular follow-up, possibly due to patient compliance issues or clinical scheduling limitations.
@@ 4.4 Justification for dataset size
Data export was manual and, therefore, time-consuming. In addition, manually annotating thousands of image pairs by two retinal experts was also time-consuming. This was the reason for not collecting a larger dataset. However, this remains larger than a similar study published by one of the organizers [Quellec2019], which involved 70 patients, as opposed to 136 here.